# Causal Explanations for Image Classifiers

### v2

@@ 6.3 Discussion
Figure 11: Cumulative plots of explanations sizes with 333 different models over all datasets. rex consistently produces smaller explanations, meaning it identifies precisely those pixels required for the classification, with fewer unnecessary pixels.
lime, when used for image classification, works best when provided with a segmentation mask of the image. For common datasets such as the ones we investigate, many high quality algorithms exist. rex does not require such existing segmentation masks. (?) conducted an investigation on brain MRI explainability using a different model and different data set and a similar, though not overlapping, set of explainability tools. On this dataset, lime’s performance is hampered by medically meaningless segmentation. They found that rex did not suffer from this problem, but performed well compared to the other tools on this different model and data.
@@
Figure 12: Photo Bombing images and output from rex
We evaluated the results produced by rex with those produced by other explainability tools on a number of standard measures proposed in the literature, such as the size of explanations, insertion and deletion curves. For the datasets containing the ground truth or some part of it, such as the segmentation information or the partial occlusion, we also measured the percentage of the explanation that is clearly outside of an area where it should be contained (such as the occlusion in partially occluded areas). For all other measures, rex shows superior performance, which answers RQ5.
@@
7 Conclusions and Future Work
@@
In this paper we described an approach to explainability of image classifiers that is rooted in actual causality and computes explanations based on the formal definitions of sufficient explanations. As the exact computation is intractable, we described a modular algorithm for computing approximate explanations using a causal ranking of parts of the input image, viewing the classifier as a black box. Our experiments demonstrate that an implementation of our algorithm in the tool rex produces results that are superior to those of state-of-the-art tools according to standard measures.
@@
In the future, we will apply rex to other modalities that lend themselves to occlusion-based reasoning, such as tabular data, time series, spectroscopy, etc. We will also explore different domains of application, where precise explanations are instrumental, in particular mission-critical and safety-critical domains, such as the healthcare AI and autonomous vehicles. The healthcare domain will require some adaptations of our algorithms, since the quality of medical images is different from the quality of general images (e.g., there are fewer colors in an X-ray than in a general image). In the automotive domain, we will adapt rex to analyze object detectors and also devise new algorithms that work in (near) real-time, to address the needs of explainability for autonomous vehicles.

### v3

@@ 6.3. Discussion
Figure 12. Cumulative plots of explanations sizes with $3$ different models over all datasets. rex consistently produces smaller explanations, meaning it identifies precisely those pixels required for the classification, with fewer unnecessary pixels.
lime, when used for image classification, works best when provided with a segmentation mask of the image. For common datasets such as the ones we investigate, many high quality algorithms exist. rex does not require such existing segmentation masks. blake2024explainable conducted an investigation on brain MRI explainability using a different model and different data set and a similar, though not overlapping, set of explainability tools. On this dataset, lime’s performance is hampered by medically meaningless segmentation. They found that rex did not suffer from this problem, but performed well compared to the other tools on this different model and data.
@@
Figure 13. Photo Bombing images and responsibility maps from rex
We evaluated the results produced by rex against those produced by other explainability tools on a number of standard measures proposed in the literature, such as the size of explanations, insertion and deletion curves. For the datasets containing the ground truth or some part of it, such as the segmentation information or the partial occlusion, we also measured the percentage of the explanation that is clearly outside of an area where it should be contained, such as the occlusion in partially occluded areas (Figure 13). For all other measures, rex shows superior performance, which answers RQ5.
@@
7. Conclusions and Future Work
@@
In this paper we described an approach to explainability of image classifiers that is rooted in actual causality and computes explanations based on the formal definitions of sufficient explanations. As the exact computation is intractable, we described a modular algorithm for computing approximate explanations using a causal ranking of parts of the input image, viewing the classifier as black-box. Our experiments demonstrate that an implementation of our algorithm in the tool rex produces results that are superior to those of state-of-the-art tools according to standard measures.
@@
In the future, we will apply rex to other modalities that lend themselves to occlusion-based reasoning, such as tabular data, time series, and spectroscopy. We will also explore different domains of application, where precise explanations are instrumental, in particular mission-critical and safety-critical domains, such as in healthcare AI and autonomous vehicles. The healthcare domain will require some adaptations of our algorithms, since the quality of medical images is different from the quality of general images (e.g., there are fewer colors in an X-ray than in a general image). In the automotive domain, we will adapt rex to analyze object detectors and also devise new algorithms that work in (near) real-time, to address the needs of explainability for autonomous vehicles.
Appendix A Our Definitions in the Actual Causality Landscape
For the ease of presentation, we first restate our definition of explanation. See 4.1
The following is the latest definition of actual cause due to Halpern (Hal15) (see (HP01b) for the original definition).
Definition A.1.
[Actual cause (Hal15)] $\vec{X}=\vec{x}$ is an actual cause of $\varphi$ in $(M,\vec{u})$ if the following three conditions hold:
AC1.:
$(M,\vec{u})\models(\vec{X}=\vec{x})$ and $(M,\vec{u})\models\varphi$.
AC2.:
AC3.:
The following lemma states their equivalence in our setting.
Lemma A.2.
Given a image classifier $\mathcal{N}$, an input $x$, and a causal model $M$ derived from them as described in Section 4, a subset $\vec{V}_{exp}$ of $\vec{V}$ is a single-context explanation of $\mathcal{N}(x)$ according to Definition 4.1 iff $\vec{V}_{exp}=0$ (that is, all variables in $\vec{V}_{exp}$ have the value 0) is an actual cause of $O=0$ in $(M,\vec{u}_{0})$ according to Definition A.1, under the assumption that a fully masked image has a different classification than $x$.
Proof.
First we observe that AC1 simply states that in the current context $\vec{u}_{0}$, the value of all variables in $\vec{V}$ is 0, and the classification of the fully masked image is different from $\mathcal{N}(x)$. Assume that EXIM1 holds. Then, assigning $\vec{V}_{exp}$ to $1$ results in restoring the original classification $\mathcal{N}(x)$, and hence $O=1$, hence AC2 holds with $\vec{W}=\emptyset$. For the other direction, assume that AC2 holds. Since the variables in $\vec{V}$ are causally independent, changing $\vec{V}_{exp}$ to $1$ does not change any other input variable, and hence we can take $\vec{W}=\emptyset$. Then, AC2 implies EXIM1. Finally, AC3 is simply the minimality condition, and is equivalent to EXIM2. ∎
We note that the theory of actual causality includes the notion of explanations as well (HP01a), but these are defined over a set of contexts, rather than a single context, so are not suitable for our purposes. We can, in general, imagine a setting in which we are searching for a subset of pixels of the input image that is sufficient to result in the original classification, regardless of the values of the rest of the pixels—this would be similar to the setting explored in Anchors (anchors). However, such a setting leads to significantly less intuitive results for explaining image classification, as also evident from the results of applying Anchors to images.
Let us now place our definition of sufficient responsibility into the actual causality landscape. For the ease of presentation, we first restate the definition. See 4.2
In the literature, the degree of responsibility is defined by Chockler and Halpern (CH04) as the quantification of actual causality, hence its precise definition is tied to the definition of the actual cause. The latest version is as follows.
Definition A.3.
[Degree of responsibility (Hal19)] For $\vec{X}=\vec{x}$ being an actual cause of $\varphi$ in $(M,\vec{u})$, the degree of responsibility of $\vec{X}=\vec{x}$ for the value of $\varphi$ in $(M,\vec{u})$ is defined as
$dr(\vec{X}=\vec{x},\varphi,M,\vec{u})=\frac{1}{|\vec{X}|+|\vec{W}|},$
for the smallest $|\vec{X}|+|\vec{W}|$ that satisfy Definition A.1.
The following is immediate from the definitions, given that in our case $\vec{W}=\emptyset$.
Proposition A.4.
Given a image classifier $\mathcal{N}$, an input $x$, a causal model $M$ derived from them as described in Section 4, and a single-context explanation $\vec{V}_{exp}\subseteq\vec{V}$ of $\mathcal{N}(x)$ according to Definition 4.1, the degree of sufficient responsibility of $\vec{V}_{exp}$ for $\varphi$ in $M$ is the degree of responsibility of $O=0$ in $(M,\vec{u}_{0})$ according to Definition A.3, under the assumption that a fully masked image has a different classification than $x$.
Indeed, the degree of sufficient responsibility of $\vec{V}_{exp}$ is $\frac{1}{\vec{V}_{exp}}$, which is exactly the degree of responsibility according to Definition A.3 for $(M,\vec{u}_{0})$, given that in our model $M$ there is no dependency between the variables.
Appendix B The Complexity of Our Definitions
In order to analyze the complexity of our definition of single-context explanation, we first need to introduce a complexity class. Papadimitriou and Yannakakis (PY84) introduced the complexity class $DP$, which consists of all languages $L$ such that there exists a language $L_{1}$ in $NP$ and a language $L_{2}$ in co-$NP$ such that $L=L_{1}\cap L_{2}$. Papadimitriou and Yannakakis showed that a number of problems of interest were $DP$ complete.
For the complexity discussion below, we use the classifier $\mathcal{N}$ as an oracle, not taking its complexity into account. Instead, we assume that the value of $O$ is computed in polynomial time using the values of variables in $\vec{V}$. Since all variables in our causal models are binary, the computation is simply a Boolean formula.
Halpern proved that the decision problem of actual causality as per Definition A.1 is $DP$-complete, and the result holds for binary models as well (Hal15). The proof, however, relies on causal models having causal dependencies between their variables. As we show below, our definition, while still intractable, is in a lower complexity class. But first we observe that for singletons, our definition is computable in polynomial time.
Proposition B.1.
If $\vec{V}_{exp}$ is a single variable $V$, the complexity of deciding whether it satisfies Definition 4.1 is polynomial in the size of the input.
Proof.
The proof is straightforward by observing that EXIM1 is checkable in polynomial time, and EXIM2 holds trivially, as singletons have no subsets. ∎
We note that the decision problem of actual cause is $NP$-complete for singletons (Hal15), indicating that our setting significantly simplifies the problem. The reduction in complexity also holds for the general case, as proved in the following theorem.
Theorem B.2.
The decision problem of explanations in Definition 4.1 is co-$NP$-complete.
Proof.
For the membership in co-$NP$, we prove that the complementary problem, of showing that a given $\vec{V}_{exp}$ is not an explanation, is in $NP$. Given a set $\vec{V}_{exp}$, checking whether it satisfies EXIM1 is polynomial. For refuting EXIM2, it suffices to find one witness subset $\vec{V}^{\prime}\subset\vec{V}$ that satisfies EXIM1, hence the complementary problem is indeed in $NP$.
For the co-NP-hardness, we show a reduction from the classic co-$NP$-complete problem $UNSAT=\{$ unsatisfiable propositional Boolean formulae }\}. Given a propositional Boolean formula $\varphi$ over the set of variables $X_{1},\ldots,X_{n}$, we construct a depth-$2$ binary causal model $M_{\varphi}$ as follows. The set of endogenous variables $\mathcal{}V$} is the set $X_{1},\ldots,X_{n},O$, where the values of $X_{1},\ldots,X_{n}$ are determined directly by the exogenous variables, and
$O=(\varphi\vee\ \bigwedge_{i=1}^{n}X_{i})\wedge(\bigwedge_{i=1}^{n}X_{i}\rightarrow\neg{\varphi}).$
Essentially, the value of $O$ is an exclusive OR of the value of $\varphi$ and the conjunction of all variables of $\varphi$. Then, $\varphi\in UNSAT$ iff the set $\vec{X}=\{X_{1},\ldots X_{n}\}$ is a single-context explanation for $O=0$ in $(M,\vec{u}_{0})$.
First, we claim that $O$ has the value $1$ only in the assignment that sets all variables in $\vec{X}$ to $1$ iff $\varphi$ is unsatisfiable. Indeed, if $\varphi$ is unsatisfiable, then its value is 0 under all assignments, and the value of the conjunction is $1$ iff all variables in $\vec{X}$ are set to $1$. For the other direction, if the value of $O$ is $1$ under the assignment that sets all variables in $\vec{X}$ to $1$, then $\varphi$ is falsified under this assignment, and since this is the only assignment that results in $O=1$, $\varphi$ is falsified under all other assignments to $\vec{X}$ as well, hence $\varphi$ is unsatisfiable.
We are now ready to prove the reduction. Assume first that $\varphi\in UNSAT$. Then, by the previous discussion, $(M,\vec{u}_{0})\models[\vec{X}\leftarrow 1](O=1)$, satisfying EXIM1, and $\vec{X}$ is the only such subset of variables, satisfying EXIM2. Therefore, $\vec{X}$ is a single-context explanation to $O$ in $M_{\varphi}$, as required.
For the other direction, assume that $\vec{X}$ is a single-context explanation to $O$ in $M_{\varphi}$. Then, $(M,\vec{u}_{0})\models[\vec{X}\leftarrow 1](O=1)$, and no subset of $\vec{X}$ satisfies this condition, hence by the previous discussion, $\varphi\in UNSAT$, completing the proof. ∎
There is, therefore, little hope to find an efficient algorithm for computing exact explanations for image classification, and the answer to RQ2 wrt precise explanations is ‘no’, as computing explanations is intractable.
While this is not directly relevant to this paper, the following observation is a direct corollary from Lemma A.2 and Theorem B.2.
Observation B.3.
In depth-2 binary causal models, the decision problem of actual causality is co-$NP$-complete.
We now turn to examining the complexity of the degree of sufficient responsibility. First, we need to introduce a new complexity hierarchy, specifically, the hierarchy of functional problems. This is because the output of the degree of the responsibility computation is a value, and not an accept/reject decision. For a complexity class $A$, the class $\mbox{FP}^{{\rm A}[\log{n}]}$ consists of all functions that can be computed by a polynomial-time Turing machine with an oracle for a problem in $A$, which on input $x$ asks a total of $O(\log{|x|})$ queries (cf. (Pap84)). A function $f(x)$ is $\mbox{FP}^{{\rm A}[\log{n}]}$ hard iff for every function $g(x)$ in $\mbox{FP}^{{\rm A}[\log{n}]}$ there exist polynomially computable functions $R,S:\Sigma^{*}\rightarrow\Sigma^{*}$ (where $\Sigma$ is the common alphabet) such that $g(x)=S(f(R(x)))$. A function $f(x)$ is complete in $\mbox{FP}^{{\rm A}[\log{n}]}$ iff it is in $\mbox{FP}^{{\rm A}[\log{n}]}$ and is $\mbox{FP}^{{\rm A}[\log{n}]}$-hard.
We prove the following result.
Theorem B.4.
The problem of computing the degree of sufficient responsibility is $\mbox{FP}^{{\rm NP}[\log{n}]}$-complete.
Proof.
The proof is very similar to the proof of complexity of the degree of responsibility in (CH04), so we present only the key points here.
For the membership in $\mbox{FP}^{{\rm NP}[\log{n}]}$, we observe that the polynomial time algorithm for computing the degree of responsibility can perform a binary search on the size of $\vec{V}_{exp}$, in each step querying an oracle for the existence of a single-context explanation $\vec{V}_{exp}$ of size $k$. By  Theorem B.2, the decision problem of the existence of explanation is co-$NP$-complete, and hence can be decided by an $NP$-oracle.
For hardness in $\mbox{FP}^{{\rm NP}[\log{n}]}$, the reduction is from the $\mbox{FP}^{{\rm NP}[\log{n}]}$-complete function problem $MINSAT$, defined in (CH04) as $MINSAT(\varphi)=k$, where $k$ is the minimal number of $1$’s in a satisfying assignment for $\varphi$. The reduction is the same as in (CH04). ∎
Theorem B.4 shows that computing the degree of sufficient responsibility in our setting is intractable, hence the need for approximation algorithms.